OpenStudy (anonymous):

8t-(-2)-14-(-5t)+53-9t

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

recall the definition of subtraction: $$\Huge a - b = a + (-b)$$

OpenStudy (anonymous):

I need help all the way around to be honest...I suck at this

OpenStudy (skullpatrol):

Once you understand: to subtract b add the opposite of b, you will be fine :)

OpenStudy (gorv):

just add the variables with t and without variable

OpenStudy (anonymous):

So I will subtract 8t from -2?

OpenStudy (anonymous):

22t and 41?

OpenStudy (skullpatrol):

Go through the expression: $$\Huge 8t-(-2)-14-(-5t)+53-9t$$ using the definition of subtraction and change all the subtractions to additions...

OpenStudy (skullpatrol):

$$\Huge 8t+(2)+(-14)+(5t)+53+(-9t) $$

OpenStudy (anonymous):

8t+(-2)+14+(-5)+53+9t?

OpenStudy (skullpatrol):

to subtract b add the opposite of b

OpenStudy (gorv):

like 2t-3-4t+4=2t-4t-3+4 (2t-4t)=-2t -3+4=1 so ur ans =-2t+1

OpenStudy (gorv):

do likewise in ur Q @Treday

OpenStudy (anonymous):

I just suck

OpenStudy (skullpatrol):

Watch please

OpenStudy (skullpatrol):